Create pseudocode

Base charge is $30.00 a month with $10.00 for each added phone. Any data usage over 3GB but less than 5GB a month is charged at $9.00 per GB (aggregate for all phones on plan). Any data at or above 5GB a month but less than 10 GB is charged at $12.00 GB. Anything 10 GB and over is $15.00 GB.

If a customer has been on the plan for more than 2 years but less than 4 years:

Base charge is $35.00 a month with $12.00 for each added phone. Any data usage over 3GB but less than 5GB a month is charged at $10.00 per GB (aggregate for all phones on plan). Any data at or above 5GB a month but less than 10 GB is charged at $15.00 GB. Anything 10 GB and over is $18.00 GB.

If a customer has been on the plan for 4 years or more:

Base charge is $50.00 a month with $9.00 for each added phone. Any data usage over 10 GB a month is charged at $14.00 per GB (aggregate for all phones on plan).
Answer #1

pseudocode for the computation of basecharge using tenure, number of phones and data usage.

  1. Start
  2. Read plan, number of years as integer in ” tenure”
  3. Read number of phones as integer in    “n_phones”
  4. Read usage of data in GB , into   ” usage”
  5. If(tenure<2) then

5.1. Base_charge=30+10.00*n_phones

5.2. if(usage>3 && usage<5)

     Charge=9.00*usage

5.3. else if(usage>5&& usage<10)

     Charge=12.00*usage

5.4 else

     Charge=15.00*usage

  1. Else If (tenure>2 && tenure <4) then

6.1.base_charge=35.00 +12.00*n_phones

6.2 if(usage>3&& usage<5)

        Charge=10.00*usage

   6.3.else if(usage>=5 && usage<10)

         Charge=15.00*usage

6.4 else

           Charge=18.00*usage

  1. Else
    1. Base_charge=50.00+9.00*n_phones
    2. If(usage>10)

Charge=14.00*usage

  1. Base_charge=base_charge+charge
  2. Display base_charge
  3. stop